# Pyvale: A Fast, Scalable, Open-Source 2D Digital Image Correlation (DIC) Engine Capable of Handling Gigapixel Images

### v1

@@ Abstract
Pyvale is an open-source software package that aims to become an all-in-one tool for sensor simulation, sensor uncertainty quantification, sensor placement optimization, and calibration/validation. Central to this is support for image-based sensors, with a dedicated Digital Image Correlation (DIC) module designed for both standalone use and integration within broader experimental design workflows. The design philosophy behind the DIC engine in Pyvale prioritizes a user-friendly Python interface with performant compiled code under the hood. This paper covers Pyvale’s 2D DIC engine design, implementation, metrological performance compared to other DIC codes, and the unique ability to handle gigapixel size scanning electron microscope (SEM) images. Finally, we compare runtimes between Pyvale and other open-source DIC codes and show strong computational performance across a range of image resolutions and thread counts.
@@ 9 Acknowledgments
We acknowledge support from UKRI through the Future Leaders Fellowship scheme, grant MR/Y015916/1.

### v2

@@ Abstract
Background: Digital Image Correlation (DIC) is a widely used full‑field measurement technique, but both open‑source and commercial packages often have limitations such as operating‑system restrictions, lack of support for deployment on computing clusters, and poor scalability to gigapixel‑scale images common in Scanning Electron Microscopy DIC (SEM-DIC). Objective: Pyvale is an open‑source software package designed for sensor simulation, uncertainty quantification, placement optimization, and calibration/validation. A key component of this is the development of a dedicated 2D DIC module intended for standalone use and integration within broader workflows. Methods: Pyvale provides a user‑friendly Python interface with performant compiled routines underneath. At its core is a multithreaded, reliability‑guided DIC algorithm. Its open‑source MIT license enables wide deployment, including on computing clusters and in automated pipelines. Results: Benchmarking with the publicly available 2D DIC challenge 2.0 dataset shows that Pyvale achieves metrological performance comparable to existing commercial and open-source DIC codes. It can correlate gigapixel‑scale image pairs in under 5 minutes on high-specification desktop workstations, with memory peaking at approximately 50 GB. Conclusions: Pyvale’s strong metrological foundation, coupled with its scalability for SEM‑DIC, positions it as a platform for sustained, community‑driven development. Its design and licensing provide a foundation for future improvements in open‑source DIC and integration into experimental design and validation workflows.
@@ 9 Acknowledgments
We acknowledge support from UKRI through the Future Leaders Fellowship scheme, grant MR/Y015916/1. We also acknowledge support from the UK Engineering and Physical Sciences Research Council (EPSRC), grant EP/W006839/1.